OpenStudy (anonymous):

A 5-foot ladder is leaning against a 4-foot wall. How far must the bottom of the ladder be from the base of the wall so that the top of the ladder rests on the top of the wall?

OpenStudy (johnweldon1993):

 We have a right triangle...we can use the pythagorean theorem \[\large a^2 + b^2 = c^2\] \[\large 4^2 + b^2 = 5^2\] isolate and solve for 'b' \[\large b = \sqrt{5^2 - 4^2}\]